OpenStudy (anonymous):

how many joules of heat are needed to raise the temperature of 10.0 g of aluminum from 22 degrees celcius to 55 degrees celcius, if the specific heat of aluminum is 0.90 joules

OpenStudy (anonymous):

Heat required=m*S(Tfinal-Tinitial)...S=0.90j/kg-degc,m=10/1000 Kg,Tinitial=22 deg,Tfinal=55 deg...put in this equation and get it.

OpenStudy (anonymous):

H = m s (T_f - T_I) H = 0,01kg * 0,90 * (55 - 22) H = 0,01kg * 0,90 * (33)ºc H = 0.287J